# Demystifying Misconceptions in Social Bots Research

### v2

@@ Introduction
In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies. Previously, studies primarily concentrated on Twitter/X, likely because its free API facilitated data acquisition. However, this accessibility ceased in 2023 when Twitter/X terminated free data access for researchers. [^7] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram. [^8] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
@@
Despite these setbacks, Europe’s Digital Service Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests. [^9] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs. However, these programs are significantly limited compared to previous access levels. Their opaque and stringent application review processes also leave their efficacy in question. Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduces new dynamics. Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors. The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community.
Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
A call for moral responsibility
Our analysis of the drawbacks and limitations of some recent studies brought to light the need for responsibility when discussing results about social bots. The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion. To this end, it is paramount to avoid propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars—such as those between proponents and opponents of social bots research.
The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society. As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses. Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019). Indeed, we believe that all the literature that propagates the misconceptions demystified in this paper provides misleading contexts in some way. Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021). Similarly, while making hyped or sensationalist claims can help publish a paper and accrue citations, such claims also contribute to creating misleading contexts and unrealistic expectations, as well as to increasing scientific tribalism.

### v3

@@ 4 Challenges in the post-API era
In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies. In the API era, most social bot studies have focused on Twitter/X due to its free API and the ease of large-scale data acquisition. This hyperfocus on Twitter/X has led to the neglect of other platforms, creating a bias in social bots research that largely ignores platforms beyond Twitter/X. However, this data policy ceased in 2023 when Twitter/X terminated free data access for researchers. [^9] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram. [^10] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
@@
Despite these setbacks, Europe’s Digital Services Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests. [^11] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs. However, these programs are significantly limited compared to previous access levels. Their opaque and stringent application review processes also leave their efficacy in question (Jaursch et al., 2024). Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduce new dynamics. Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors. The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community. Moreover, the limited number of participants, which has not yet reached a critical mass, and the difficulty in creating a reliable ground truth further contributes to hindering bot detection research on these new platforms. Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
5 A call for moral responsibility
Our analysis of the drawbacks and limitations of recent studies brought to light the need for responsibility when discussing results about social bots. The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion. To this end, it is paramount to avoid repeating the mistakes and propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars.
The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society. As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses. Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019). Perpetuating the methodological and conceptual issues addressed in this paper contributes to creating flawed or unreliable research. Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021). Similarly, while making hyped or sensationalist claims can help publish a paper and accrue citations, such claims also contribute to creating misleading contexts and unrealistic expectations, as well as to increasing scientific tribalism.